From GlueXWiki
Jump to: navigation, search

September 25, 2014 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. DAQ (Beni)
  2. CDC update (Mike)
  3. FDC update (Vlad, Lubomir)
  4. Electronics (Fernando)
  5. Engineering
  6. Commissioning Wiki Page (all)
  7. Other


Participants: Curtis, Naomi (CMU), Mike, Eugene, Dave, Chris, Nick, Simon, Beni, Bill, Fernando, Cody, Vlad, and Lubomir (JLab).


- Beni: CDC+FDC DAQ modules have been moved to the top trigger crate where the GTP is. Sasha is ready with the BCAL trigger, but there's more work to be able to read the CDC+FDC+BCAL. The plan is to use again scintillator trigger over the weekend without moving the trigger modules back to the bottom crate. Beni is working on interface between the data base parameters and the Sasha's initialization files.

- Lubomir looked at the runs with mode 8, where you have both raw and processed data. In the figures linked above we compare the processed data with the one obtained from raw data, by simulating the firmware offline. There's agreement for the pedestal calculation, but not for the peak value and the timing. Later it turned out f125 is reporting up to two peaks, while offline is looking at the first peak only. So far it is not clear which are the two peaks in their order, since the max number of peaks was set to 4. Cody will look into this. Another problems is that for certain channels (most in ROCFDC2) we see the pedestals equal to zero. It turns out this is a feature of the firmware when the pedestal is bigger than the peak value.

CDC update

- Mike: 2-3 3M connectors should be fixed. Other than that a few channels have problems. Single channel HV modules Bertan type, with sensitive current reading, will be used for 2-3 channels to help the beam alignment. They will have to be turned on before the start of the beam alignment and will read the current with a camera.

FDC update

- Vald reported, we don't see signal from one wire. After the meeting Nick and Lubomir checked the signals with a pulser (applied on the pre-amp input): there were signals visible, but the signal on this channel weaker than the rest. Most likely the pre-amp connector doesn't seat well. Vlad is analyzing different runs to figure out when the channels stopped working.


- Fernando: ethernet cable temporary connected to FDC MPOD, many ethernet cables without labels.

- Cody: the first article f125 will be here in mid-October. The rest November December.


- Bill: The chiller will be send in January to mid-March for fixing to the company. They will replace the compressor, fix the low level switch, and test it, all for free. Eugene: we will need it mid-March for the next run.